OpenStudy (anonymous):

A box weighing 18 N requires a force of 6.0 N to drag it at a constant rate. What is the coefficient of sliding friction? 0.33 3.0 6.0 18

OpenStudy (anonymous):

@Abhisar

OpenStudy (abhisar):

Frictional force = \(\sf mg \mu\) = 6N Find the value of \(\sf \mu\) by setting up an equation.

OpenStudy (abhisar):

mg= weight

OpenStudy (abhisar):

getting it?

OpenStudy (anonymous):

6.0?

OpenStudy (abhisar):

no, 18*\(\sf \mu\)=6, can u find \(\sf \mu\) now?

OpenStudy (anonymous):

nope

OpenStudy (abhisar):

y? \(\sf 18 \times \mu = 6\\ \mu = 6/18 =?\)

OpenStudy (anonymous):

3.0?

OpenStudy (abhisar):

you are dividing it wrong!

OpenStudy (abhisar):

it's \(\frac {6}{18}~and~not~\frac{18}{6}\)

OpenStudy (anonymous):

oh ok thnx